Based on the table, the interest rate on a 3-month Eurodollar deposit made in December of 2019 is expected to be

Dec 2019 97.31

Mar 2020 97.13

Jun 2020 96.97

Sept 2020 96.86
Answer #1

So here in your Question you asked for interest rate, So we have following

Present Value(Discounted Deposite): 97.31

Future Value (Face Value): 100

Period: 3 months: 3/12

i : x  

So formula for future value is as follows

PV(1+i)n = FV

97.31( 1+ x )3/12 = 100

( 1+ x )3/12 = 100 / 97.31

( 1+ x ) = (1.02764)12/4

( 1+ x ) = 1.11524

x = 0.11524

i = 11.52%

In case you want to solve in excel or on financial Calculator use following steps:

Excel: =rate(3/12,0,-97.31,100)

Financial Calculator: n: 1/12, PV: -97.31, PMT: 0, FV: 100